# GEBench: Benchmarking Image Generation Models as GUI Environments

### v1

@@ 5.1 Hierarchy of Task Difficulty: From Local Mimicry to Global Reasoning Failure
The experimental results, as shown in Table 1 and Figure 4, reveal a pronounced inverse correlation between task complexity and model performance, highlighting a fundamental deficiency in current models’ deep understanding of GUI mechanics. In single-step transition scenarios, leading models team2023gemini; openai2025gptimage achieve robust scores exceeding 80% through powerful visual synthesis. However, this “illusory prosperity" is largely driven by a shortcut mapping from instructions to visual patterns; rather than mastering the underlying interactive logic, models primarily leverage statistical distribution fitting to match expected GUI states. This structural weakness becomes evident as tasks extend to multi-step trajectories. The significant inward “shrinkage" of the radar chart demonstrates an inability to maintain temporal coherence, where the lack of explicit state-space logic leads to a severe logical disconnect when processing high-complexity interaction flows.
@@ 5.2 In-depth Bottleneck Analysis: Qualitative Insights from Failure Cases
By synthesizing qualitative evidence, as shown in Figure 7, we identify three primary technical bottlenecks that hinder reliable GUIs generation. The first is the failure of text rendering accuracy. Qualitative cases reveal that while models like Nano Banana Pro team2023gemini encounter deformations in complex layouts, open-source models frequently exhibit severe character overlapping and semantic corruption. This suggests that models treat text as a local texture rather than a symbolic unit with rigid structural information, leading to unreadable characters in layout-dense environments without hard topological constraints.
@@
Multi-dimensional analysis through GE-Score reveals a critical paradox: visual fidelity does not equate to functional viability. Models such as GPT-image-1.5 openai2025gptimage generate GUIs with exceptional composition and clarity, earning high QUAL scores. However, granular functional inspection reveals that this “visual over-optimism" is often deceptive, as these aesthetically pleasing images frequently contain hallucinated widgets or illogical layouts. This reinforces that benchmarking image generation models as GUI environments must be predicated upon a assessment of temporal coherence and interactive logic, which take precedence over general-domain visual fidelity.
@@ Appendix A Evaluation Framework
In this section, we detail the proposed evaluation framework. The evaluation framework operates through a systematic three-stage pipeline designed to rigorously benchmark image generation models as GUI environments. The process initiates with Image Generation, where image generation models are tasked with generating visual outputs across five distinct task categories: Single-step Visual Transition (single-step), Multi-step Planning (multi-step), Zero-shot Virtual GUI (fiction-app), Rare Trajectory Synthesis (real-app), and Grounding-based Generation (grounding). Subsequently, the generated samples undergo a VLM-as-a-Judge process employing a evaluator strategy. We leverage VLMs, specifically GPT-4o hurst2024gpt, Gemini-3-Pro-Native google2025gemini3 and Qwen3-vl-235b-a22b-thinking bai2025qwen3vltechnicalreport, to assess the results along five critical dimensions: Goal Achievement (GOAL), Interaction Logic (LOGIC), Consistency (CONS), UI Plausibility (UI), and Visual Quality (QUAL). Finally, the framework concludes with Metrics Analysis, where the calculated scores are validated through statistical verification, pattern analysis, and human relevance alignment to ensure robust and meaningful benchmarking results.
@@ Appendix B Detailed Performance On GEBench Using Different Judges
  • Table A1: Detailed scores assigned by Gemini-3-Pro-Native google2025gemini3 as the judge.
  • Table A2: Detailed scores assigned by GPT-4o hurst2024gpt as the judge.
  • Table A3: Detailed scores assigned by Qwen3-vl-235b-a22b-thinking bai2025qwen3vltechnicalreport as the judge.
@@ Appendix C Detailed Rubric on five tasks
Table A1: Detailed Performance on GEBench using Gemini-3-Pro-Native google2025gemini3 as Judge
@@
Table A2: Detailed Performance on GEBench using GPT-4o hurst2024gpt as Judge
@@
Table A3: Detailed Performance on GEBench using Qwen3-vl-235b-a22b-thinking bai2025qwen3vltechnicalreport as Judge

### v2

@@ 5.1 Hierarchy of Task Difficulty: From Local Mimicry to Global Reasoning Failure
The experimental results, as shown in Table 1 and Figure 4, reveal a pronounced inverse correlation between task complexity and model performance, highlighting a fundamental deficiency in current models’ deep understanding of GUI mechanics. In single-step transition scenarios, leading models (team2023gemini; openai2025gptimage) achieve robust scores exceeding 80% through powerful visual synthesis. However, this “illusory prosperity" is largely driven by a shortcut mapping from instructions to visual patterns; rather than mastering the underlying interactive logic, models primarily leverage statistical distribution fitting to match expected GUI states. This structural weakness becomes evident as tasks extend to multi-step trajectories. The significant inward “shrinkage" of the radar chart demonstrates an inability to maintain temporal coherence, where the lack of explicit state-space logic leads to a severe logical disconnect when processing high-complexity interaction flows.
@@ 5.2 In-depth Bottleneck Analysis: Qualitative Insights from Failure Cases
By synthesizing qualitative evidence, as shown in Figure 7, we identify three primary technical bottlenecks that hinder reliable GUIs generation. The first is the failure of text rendering accuracy. Qualitative cases reveal that while models like Nano Banana Pro (team2023gemini) encounter deformations in complex layouts, open-source models frequently exhibit severe character overlapping and semantic corruption. This suggests that models treat text as a local texture rather than a symbolic unit with rigid structural information, leading to unreadable characters in layout-dense environments without hard topological constraints.
@@
Multi-dimensional analysis through GE-Score reveals a critical paradox: visual fidelity does not equate to functional viability. Models such as GPT-image-1.5 (openai2025gptimage) generate GUIs with exceptional composition and clarity, earning high QUAL scores. However, granular functional inspection reveals that this “visual over-optimism" is often deceptive, as these aesthetically pleasing images frequently contain hallucinated widgets or illogical layouts. This reinforces that benchmarking image generation models as GUI environments must be predicated upon a assessment of temporal coherence and interactive logic, which take precedence over general-domain visual fidelity.
@@ Appendix A Evaluation Framework
In this section, we detail the proposed evaluation framework. The evaluation framework operates through a systematic three-stage pipeline designed to rigorously benchmark image generation models as GUI environments. The process initiates with Image Generation, where image generation models are tasked with generating visual outputs across five distinct task categories: Single-step Visual Transition (single-step), Multi-step Planning (multi-step), Zero-shot Virtual GUI (fiction-app), Rare Trajectory Synthesis (real-app), and Grounding-based Generation (grounding). Subsequently, the generated samples undergo a VLM-as-a-Judge process employing a evaluator strategy. We leverage VLMs, specifically GPT-4o [hurst2024gpt], Gemini-3-Pro-Native [google2025gemini3] and Qwen3-vl-235b-a22b-thinking [bai2025qwen3vltechnicalreport], to assess the results along five critical dimensions: Goal Achievement (GOAL), Interaction Logic (LOGIC), Consistency (CONS), UI Plausibility (UI), and Visual Quality (QUAL). Finally, the framework concludes with Metrics Analysis, where the calculated scores are validated through statistical verification, pattern analysis, and human relevance alignment to ensure robust and meaningful benchmarking results.
@@ Appendix B Detailed Performance On GEBench Using Different Judges
  • Table A1: Detailed scores assigned by Gemini-3-Pro-Native [google2025gemini3] as the judge.
  • Table A2: Detailed scores assigned by GPT-4o [hurst2024gpt] as the judge.
  • Table A3: Detailed scores assigned by Qwen3-vl-235b-a22b-thinking [bai2025qwen3vltechnicalreport] as the judge.
@@ Appendix C Detailed Rubric on five tasks
Table A1: Detailed Performance on GEBench using Gemini-3-Pro-Native [google2025gemini3] as Judge
@@
Table A2: Detailed Performance on GEBench using GPT-4o [hurst2024gpt] as Judge
@@
Table A3: Detailed Performance on GEBench using Qwen3-vl-235b-a22b-thinking [bai2025qwen3vltechnicalreport] as Judge